OpenStudy (anonymous):

Hello! I have an issue... Here is the question: Expand (2+x)^4 and simplify your result. Hence, find the term in x^2 in (2+x))^4*(1+(1/x^2))

OpenStudy (freckles):

(2+x)^2=(2+x)(2+x)=2(2+x)+x(2+x)=2(2)+2(x)+x(2)+x(x)=4+2x+2x+x^2 (2+x)^2=4+4x+x^2 (2+x)^4=(2+x)^2*(2+x)^2 (2+x)^4=(4+4x+x^2)(4+4x+x^2) just multiply as I did above